OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = one halfx - 8. y = one halfx + 7 y = one halfx + 5 y = 2x + 7 y = 2x + 5

OpenStudy (larseighner):

What is true of parallel lines on a graph? That is, how would know lines are parallel?

OpenStudy (anonymous):

I really don't know any of this. Can you help me work out the equation step by step?

OpenStudy (larseighner):

Parallel lines on a graph have the same slope. As it happens the line you are asked about and all of the choices are expressed in slope-intercept form. What is the slope of the line you are asked about?

OpenStudy (triciaal):

the equation of a line in slope intercept form is y= mx + b where (x, y) is any point on the line m is the slope of the line and b the intercept. You are given the line y= (1/2)x -8. equating this with y = mx + b what is the slope of this line?

OpenStudy (anonymous):

1/2 is the slope right @triciaal

OpenStudy (larseighner):

That's right. Now which answerS have the same slope?

OpenStudy (anonymous):

@LarsEighner A and B

OpenStudy (larseighner):

Okay. Now it is narrowed down. Which of these two contains the point (-2,6)?

OpenStudy (anonymous):

I don't know. Can you help me with this part?

OpenStudy (larseighner):

Okay (-2,6) is the point at which x= -2 and y=6. If you substitute in the equations, which of them is true?

OpenStudy (anonymous):

Dude I don't know. Tell me how to do it, I can figure it out.

OpenStudy (larseighner):

If you substite y = 6 and x= -2 in B, you get 6 = 1/2(-2) + 5 6 = -1 + 5 Is that a true statement?

OpenStudy (anonymous):

So it's B? Oh so you just plug in the numbers for x and y

OpenStudy (larseighner):

Well, no. When you plug into B you get 6= 1/2(-2) + 5 6= -1 + 5 6= 4 Is that true?

OpenStudy (anonymous):

I don't know

OpenStudy (triciaal):

parallel lines have the same slope as stated by @LarsEighner y = mx + b line passing through the point (-2, 6) m = 1/2 you have x, y and m so now we can find b what is b?

OpenStudy (larseighner):

You really do not know if 6=4? How about we trade my $4 for your $6. Would you like a trade like that?

OpenStudy (anonymous):

wait give me a minute.

OpenStudy (triciaal):

y = mx + b 6 = 1/2*(-2) + b 6= -1 + b b =

OpenStudy (anonymous):

Alright, so it's a because when you plug in the numbers you get 6=1/2-1+7 6=-1+7 6=6

OpenStudy (triciaal):

write the equation of the line for any point (x, y) slope 1/2 and b =7

OpenStudy (triciaal):

note that if you were not given the choices then this is the way to work through the problem

OpenStudy (larseighner):

Right. @triciaal 's approach will work too. Look by knowing what slope is, you could double your guessing chances on an item like this. By plugging into only the candidate answers, you can find the right answer with minimal math. Half of math is not doing stuff you don't have to do to get the answer.